Calculate the normal and the geodesic curvatures of the following curves on
the given surfaces:
(a) The circle γ(t) = (cost, sin t, 1) on the paraboloid σ(u, v) = (u, v, u^2 + v^2).
(b) The right circular helix γ(θ) = (a cos θ, a sin θ, bθ) on the right circular cylinder
σ(u, v) = (a cos u, a sin u, v), where a, b > 0 are constants.
1
Expert's answer
2020-11-12T19:26:30-0500

The unit normal vector is given by (http://mathonline.wikidot.com/the-frenet-serret-formulas): "N(s)=\frac{T'(s)}{||T'(s)||}" , where "T(s)=\frac{\vec{r}'(s)}{||\vec{r}'(s)||},"

where "r(s)" is an arc-length parametrization of "\vec{r}(t)".

(a) We point out that one has an arc-length parametrization in this case. The length of the circle is "2\pi" and it corresponds to the interval "[0,2\pi)" for possible values of "t". Thus, t=s. We have:

"r'(t)=(-sin(t),cos(t),0)\Rightarrow ||\vec{r}'(t)||=1."

"T(t)=(-sin(t),cos(t),0)"; "T'(t)=(-cos(t),-sin(t),0)";

"N(t)=(-cos(t),sin(t),0)".

The geodesic curvature is (https://encyclopediaofmath.org/wiki/Geodesic_curvature):

"k_g=\frac{(r',r'',n)}{|r'|^3}" , where "n" is a normal vector for the surrface. The normal vector is:

"n=(2u,2v,-1)" (see https://mathworld.wolfram.com/NormalVector.html).

After substitution of the curve equation, we get that "n=(2cos\,t,2sin\,t,-1)" on the points of the curve. Thus, "k_g=\begin{vmatrix}
 -sin\,t & cos\,t &0 \\
 -cos\,t & -sin\,t &0\\
2cos\,t & 2sin\,t &-1
\end{vmatrix}=-1"

(b) We calculate the length of the curve between "[0,2\pi)" and get:

"I=\int_{0}^{2\pi}\sqrt{(x'(\theta))^2+(y'(\theta))^2+(z'(\theta))^2}d\theta=\int_{0}^{2\pi}\sqrt{a^2sin^2(\theta)+a^2cos^2(\theta)+b^2}d\theta=2\pi\sqrt{a^2+b^2}"

We make a change of variables: "s=\theta\sqrt{a^2+b^2}" . Then the curve takes the form:

"x(s)=a\,cos(\frac{s}{\sqrt{a^2+b^2}}); y(s)=a\,sin(\frac{s}{\sqrt{a^2+b^2}});\,\,z(s)=b\frac{s}{\sqrt{a^2+b^2}}" . In case we consider an interval "s\in[0,2\pi\sqrt{a^2+b^2})" we receive an arc-length parametrization.

"\vec{r}'(s)=(-\frac{a}{\sqrt{a^2+b^2}}\,sin(\frac{s}{\sqrt{a^2+b^2}}),\frac{a}{\sqrt{a^2+b^2}}\,cos(\frac{s}{\sqrt{a^2+b^2}}),\frac{b}{\sqrt{a^2+b^2}})" ; "||\vec{r}'(s)||=1".

"T(s)=\vec{r}'(s)" ; "T'(s)=(-\frac{a}{{a^2+b^2}}\,sin(\frac{s}{\sqrt{a^2+b^2}}),\frac{a}{{a^2+b^2}}\,cos(\frac{s}{{a^2+b^2}}),0)" ;

"N'(s)=\frac{T'(s)}{||T'(s)||}=\frac{a^2+b^2}{|a|}(-\frac{a}{{a^2+b^2}}\,sin(\frac{s}{\sqrt{a^2+b^2}}),\frac{a}{{a^2+b^2}}\,cos(\frac{s}{{a^2+b^2}}),0)"

The geodesic curvature is: "k_g=\frac{(r',r'',n)}{|r'|^3}" .The normal vector for the parametric surface is: "n=r_u\times r_v=\begin{vmatrix}
 i & j& k \\
 -a\,sin\,u & a\,cos\,u&0\\
 0 & 0&1
\end{vmatrix}=(a\,cos\,u,a\,sin\,u,0)" . We set "u=\theta" and get the following curvature:

"k_g=\frac{(r',r'',n)}{|r'|^3}=\frac{\begin{vmatrix}
 -a\,sin\,\theta & a\,cos\,\theta\,&b \\
 -a\,cos\,\theta & -a\,sin\,\theta&0\\
a\,cos\,\theta & a\,sin\,\theta&0
\end{vmatrix}}{(a^2+b^2)^{3/2}}=\frac{1}{(a^2+b^2)^{3/2}}b\,\,0=0"